Lincoln High School's basketball team won the regional playoffs scoring a total of 60 points no including free throws. The team made a total of 26 baskets some were 2-points shots and the rest were 3-points shots. How many 2-point shots did the team make?
Answer by greenestamps(13200)



Algebraically....

x = # of 2-point shots
y = # of 3-point shots

x+y = 26 (26 total shots)
2x+3y = 60 (60 total points)

Solve the pair of equations by whatever method you choose.

If an algebraic solution is not required, here is an easy way to find the answer using logical reasoning and simple mental arithmetic.

(1) If all 26 shots had been 2-point shots, the total score would have been 52, which is 8 points short of the actual total.
(2) Each 3-point shot counts 1 point more than each 2-point shot.
(3) Therefore the number of 3-point shots must have been 8.
(4) And that means the number of 2-point shots was 26-8 = 18.